Assume that females have pulse rates that are normally distributed with a mean of μ=76.0 beats per minute and a standard deviation of σ=12.5 beats per minute. a. If 1 adult female is randomly selected, find the probability that her pulse rate is between 69 beats per minute and 83 beats per minute. The probability is (Round to four decimal places as needed.)


1
Expert's answer
2021-08-02T09:35:56-0400

Solution:

We are given that females have pulse rates that are normally distributed with


"\mu=76, \sigma=12.5.""z={{\overline{x}-\mu} \over {\sigma / \sqrt{n}}}""z={{\overline{x}-76} \over {12.5 / \sqrt{1}}}={{\overline{x}-76} \over {12.5 }}""P(69<\overline{x}<83)=P(\overline{x}<83)-P(\overline{x}<69)""=P(z<{{83-76} \over {12.5 }})-P(z<{{69-76} \over {12.5 }})
\\=P(z<0.56)-P(z<-0.56)=P(z<0.56)-[1-P(z\le0.56)]
\\=2\times0.71226-1
\\=0.42452"